OpenStudy (anonymous):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius

OpenStudy (anonymous):

@experimentX

OpenStudy (nurali):

3000 = (h * l)+ pi * r^2 l = 2*pi*r 3000 = (h * 2*pi*r)+ pi * r^2 3000 = (50 * 2*pi*r)+ pi * r^2 3000 = 50pir + pir^2 http://www.wolframalpha.com/input/?i=3000+%3D+50pir+%2B+pir%5E2

OpenStudy (anonymous):

ty @Nurali :)